OpenStudy (anonymous):

wo drivers moving in opposite directions on an 18-mile circular racetrack cross each other at a certain point. If one of them has an average speed of 1.2 miles per minute and the other has an average speed of 1.8 miles per minute, how many minutes will it take for them to cross each other again? a) 3 b)4 c)6 d)9 e)12

OpenStudy (anonymous):

Two*

OpenStudy (anonymous):

@Michele_Laino can you help me out in this one ?

OpenStudy (anonymous):

i got it :) @Michele_Laino

OpenStudy (michele_laino):

when they cross for first time, they traveled the distance \( \large d_1, d_2 \) such that: \[\Large {d_1} + {d_2} = 18\] where: \[\Large \begin{gathered} {d_1} = {v_1}{t_0} = 1.2 \cdot {t_0} \hfill \\ {d_2} = {v_2}{t_0} = 1.8 \cdot {t_0} \hfill \\ \end{gathered} \] \( \large t_0\) is the time at they cross each other for first time. Afetr a substitution, I get: \[\Large 1.2 \cdot {t_0} + 1.8 \cdot {t_0} = 18\] Please solve that equation for \( \large t_0 \)